# h1: Bootstrapping LLMs to Reason over Longer Horizons via Reinforcement Learning

### v1

@@ Abstract
Large language models excel at short-horizon reasoning tasks, but performance drops as reasoning horizon lengths increase. Existing approaches to combat this rely on inference-time scaffolding or costly step-level supervision, neither of which scales easily. In this work, we introduce a scalable method to bootstrap long-horizon reasoning capabilities using only existing, abundant short-horizon data. Our approach synthetically composes simple problems into complex, multi-step dependency chains of arbitrary length. We train models on this data using outcome-only rewards under a curriculum that automatically increases in complexity, allowing RL training to be scaled much further without saturating. Empirically, our method generalizes remarkably well: curriculum training on composed 6th-grade level math problems (GSM8K) boosts accuracy on longer, competition-level benchmarks (GSM-Symbolic, MATH-500, AIME) by up to $2.06\times$. Importantly, our long-horizon improvements are significantly higher than baselines even at high pass@k, showing that models can learn new reasoning paths under RL. Theoretically, we show that curriculum RL with outcome rewards achieves an exponential improvement in sample complexity over full-horizon training, providing training signal comparable to dense supervision. h1 therefore introduces an efficient path towards scaling RL for long-horizon problems using only existing data.
@@ 1 Introduction
Our results in Sections 4 and 5 show that not only does this synthetic curriculum generalize to long in-domain multi-step problems, but it also transfers to harder benchmarks with longer problems such as MATH-500 and AIME that implicitly require LHR. Importantly, we show that long-horizon reasoning depends on more than just improving single step accuracy, and in Section 3 we present our method that directly targets the capabilities needed for better performance. We evaluate our long-horizon trained models versus other strong baselines up to pass@128 and show that while improvements obtained from RLVR on standard data are bounded by the base model’s capabilities (Yue et al., 2025), our method always performs significantly better. These improvements reflect new skills acquired through curriculum-based training on compositional tasks, supported by both empirical evidence and theory (Section 6). In Section 7, we further analyze compute–data tradeoffs, showing how scaling compute can substitute for scarce long-horizon data in real-world scenarios.
@@
    Empirical evidence of transfer to significantly harder benchmarks (MATH-500, AIME, GSM Symbolic, LongBench-v2, Hash-hop) while training on compositional GSM8K data.
@@ 5 Generalization to harder benchmarks
Transfer to long-context benchmarks.
We now evaluate our GSM8K LHR models on OOD long-context benchmarks to see if the state tracking capabilities ($\sigma_{j}$) from Section 3 improve. We test two main long-context benchmarks: LongBench-v2 (Bai et al., 2025) and Hash-hop (Magic, 2024). LongBench-v2 measures understanding and reasoning over QA documents, long-dialogue, repositories, etc. (with 8k–2M words). Hash-hop tests ultra-long-context storage, retrieval, and multi-hop variable tracing by making models follow shuffled chains of random hash $\rightarrow$ hash pairs. Table D summarizes our results, with a $35.00\%\rightarrow 37.90\%$ improvement on LongBench-v2 and a $15.98\%\rightarrow 18.73\%$ improvement on Hash-hop, both completely unrelated to GSM8K.
@@
This transfer patterns aligns with our pass@k capability improvement results from Section 4 and our theoretical framing. Tasks requiring sequential dependent reasoning, such as AIME or GSM-Symbolic problems, benefit from improved long-horizon reasoning capabilities that were learned on much simpler composed tasks. Crucially, improvements in aspects such as state-tracking ($\sigma_{j}$) are also observable from our long-context evals. Our results indicate that a curriculum of simple explicit-horizon tasks can bootstrap advanced reasoning, providing a scalable path where composing problems at the edge of what can be solved would push capabilities further without new annotations.
@@ 8 Discussion
While the goal of our paper is to introduce an early method for improving long-horizon reasoning, we see two promising directions for extensions. One is incorporating new sources of atomic skills beyond GSM8K. The other is creating new chaining methods that expands the serial dependency structure in our current method. We believe these two paths would offer useful extensions to the framework we introduce in this paper and further improve long-horizon reasoning capabilities.
@@ C.2 Searching for the minimum cost distribution under the same budget
Table 7: Long Horizon MATH Training on Qwen-2.5-7B Instruct. Curriculum stages lead to improvements in in-domain performance and generalization metrics. We believe these numbers will improve further with better hyperparameter tuning and more training compute. Currently, this result aims to show that our method can indeed transfer to larger models and harder training datasets.

### v2

@@ Abstract
Large language models excel at short-horizon reasoning tasks, but performance drops as reasoning horizon lengths increase. Existing approaches to combat this rely on inference-time scaffolding or costly step-level supervision, neither of which scales easily. In this work, we introduce a scalable method to bootstrap long-horizon reasoning capabilities using only existing, abundant short-horizon data. Our approach synthetically composes simple problems into complex, multi-step dependency chains of arbitrary length. We train models on this data using outcome-only rewards under a curriculum that automatically increases in complexity, allowing RL training to be scaled much further without saturating. Empirically, our method generalizes remarkably well: curriculum training on composed 6th-grade level math problems (GSM8K) boosts accuracy on longer, competition-level benchmarks (GSM-Symbolic, MATH-500, AIME) by up to $2.06\times$. It also transfers significantly to diverse out-of-distribution ReasoningGym domains and long-context benchmarks, indicating broader generalization. Importantly, our long-horizon improvements are significantly higher than baselines even at high pass@k, showing that models can learn new reasoning paths under RL. Theoretically, we show that curriculum RL with outcome rewards achieves an exponential improvement in sample complexity over full-horizon training, providing training signal comparable to dense supervision. h1 therefore introduces an efficient path towards scaling RL for long-horizon problems using only existing data.
@@ 1 Introduction
Our results in Sections 4 and 5 show that not only does this synthetic curriculum generalize to long in-domain multi-step problems, but it also transfers to harder benchmarks with longer problems such as MATH-500 and AIME that implicitly require LHR. We also observe significant generalization on diverse out-of-distribution ReasoningGym domains (Stojanovski et al., 2025) and long-context benchmarks. Importantly, we show that long-horizon reasoning depends on more than just improving single step accuracy, and in Section 3 we present our method that directly targets the capabilities needed for better generalizable performance. We evaluate our long-horizon trained models versus other strong baselines up to pass@128 and show that while improvements obtained from RLVR on standard data are bounded by the base model’s capabilities (Yue et al., 2025), our method always performs significantly better. These improvements reflect new skills acquired through curriculum-based training on compositional tasks, supported by both empirical evidence and theory (Section 6). Finally, in Section 7, we further analyze compute–data tradeoffs, showing how scaling compute can substitute for scarce long-horizon data in real-world scenarios.
@@
    Empirical evidence of transfer to significantly harder out-of-distribution benchmarks (MATH-500, AIME, GSM Symbolic, ReasoningGym, LongBench-v2, Hash-hop) while training on composed GSM8K data, outperforming other RL training baselines.
@@ 5 Generalization to harder benchmarks
Transfer to long-context benchmarks.
We now evaluate our GSM8K LHR models on OOD long-context benchmarks to see if the state tracking capabilities ($\sigma_{j}$) from Section 3 improve. We test two main long-context benchmarks: LongBench-v2 (Bai et al., 2025) and Hash-hop (Magic, 2024). LongBench-v2 measures understanding and reasoning over QA documents, long-dialogue, repositories, etc. (with 8k–2M words). Hash-hop tests ultra-long-context storage, retrieval, and multi-hop variable tracing by making models follow shuffled chains of random hash $\rightarrow$ hash pairs. Tables 3 and D summarize our results, with a $35.00\%\rightarrow 37.90\%$ improvement on LongBench-v2 and a $15.98\%\rightarrow 18.73\%$ improvement on Hash-hop, both completely unrelated to GSM8K.
Transfer to non-mathematical reasoning benchmarks.
We also test our long-horizon trained models on ReasoningGym (Stojanovski et al., 2025) domains to evaluate whether the horizon-dependent reliability improvements generalize to non-mathematical but verifiable reasoning tasks. ReasoningGym consists of a diverse set of reasoning environments that allow us to evaluate cross-domain transfer and skill generalization. Specifically, we test across logic (propositional logic), graphs (largest island), algorithmic problems (sentence reordering and matrix manipulation), arithmetic (decimal arithmetic), and geometry. These problems require working memory, graph traversal, multi-step rule following, and correct final answers. On ReasoningGym, long-horizon training on composed GSM8K significantly outperforms both the Instruct model and RLVR trained on normal GSM8K. $h1$ generalizes from $22.90\%\rightarrow 47.10\%$ on propositional logic, $15.00\%\rightarrow 22.50\%$ on graph problems (largest island), $9.60\%\rightarrow 18.80\%$ on algorithmic sentence reordering, and $2.70\%\rightarrow 4.20\%$ on algorithmic matrix manipulation. Performance on geometry drops from $3.70\%\rightarrow 2.60\%$ and on games (game of life) from $76.20\%\rightarrow 74.90\%$. Overall, skills learnt from long-horizon training generalize well to out-of-distribution reasoning problems. See Table 3.
Generalization to ReasoningGym domains Long-Context Benchmarks
Model / setting
Propositional
logic
Graphs
(largest island)
Algorithmic
(sentence reorder)
Algorithmic
LongBench-
v2
Hash-
hop
Instruct $22.90$ $15.00$ $9.60$ $2.70$ $35.00$ $15.98$
Standard RLVR $12.40$ $17.00$ $9.80$ $3.90$ $35.30$ $14.76$
Long-horizon RL 47.10 22.50 18.80 4.20 37.90 18.73
Table 3: Long-horizon training on composed GSM8K problems generalizes remarkably well to OOD ReasoningGym domains and Long-Context Benchmarks, outperforming length-1 (standard) RLVR and the Instruct model. We use default ReasoningGym configurations for our evaluations.
@@
This transfer patterns aligns with our pass@k capability improvement results from Section 4 and our theoretical framing. Tasks requiring sequential dependent reasoning, such as AIME or GSM-Symbolic problems, benefit from improved long-horizon reasoning capabilities that were learned on much simpler composed tasks. Crucially, improvements in aspects such as state-tracking are also observable from our long-context evals. Our results indicate that a curriculum of simple explicit-horizon tasks can bootstrap advanced reasoning, providing a scalable path where composing problems at the edge of what can be solved would push capabilities further without new annotations. Importantly, our method replicates on other model families (using the Llama-3.2-3B Instruct model) and additional results are presented in Appendix D.
@@ 8 Discussion
While the goal of our paper is to introduce an early method for improving long-horizon reasoning, we see two promising directions for extensions. One is incorporating new sources of atomic skills beyond GSM8K. The other is creating new chaining methods that expands the serial dependency structure in our current method. We believe these two paths would offer useful extensions to the framework we introduce in this paper and further improve long-horizon reasoning capabilities. Our codebase is open source and available at https://github.com/AlesyaIvanova/h1.
@@ C.2 Searching for the minimum cost distribution under the same budget
Table 8: Long Horizon MATH Training on Qwen-2.5-7B Instruct. Curriculum stages lead to improvements in in-domain performance and generalization metrics. We believe these results will improve further with better hyperparameter tuning and more training compute. Currently, this result aims to show that our method can indeed transfer to larger models and harder training datasets.